QUESTIONS 1-3 GO WITH THE FOLLOWING: Terry Dactal has compiled the financial information displayed below. Which of the following is Terrys net worth? $8,600 $8,450 $8,500 $14,300 $40,900 Salaries $72,400 Credit Card Balance Cash on Had 1,500 Utilities paid to date Coin Collection $2,350 Jewelry value Home value $335,000 Auto loan balance Stock Portfolio $18,500 value 1967 Ford Mustang value Grocery Expenses $7,550 Checking account $3,200 Mortgage Balance $278,600 Property Taxes owed $1,750 Mortgage loan payments made $19,500 New York vacatiorn expenses paid Income taxes paid-to-date Student loan balance $26,200 Auto loan payments $4,200 paid $6,600 Clothing/entertainment expense Insurance premiums paid $9,100 $5,000 Interest earned $400 $5,500 O $ 80,500 O $ 100,400 O $ 89,200 O $ 94,400

Which of the following is Terry's cash surplus/deficit?

$ 13,800

$6,900

$ 7,100

$10,500

If Terry had a solvency ratio of 42% the previous year, which of the following statements is correct?

there is a slight improvement from the previous year

there is a drastic improvement from the previous year

there is a slight decline from the previous year

there is a drastic decline from the previous year
Answer #1
Q-1) Net worth = Total assets-Total liabilities = (1500+2350+335000+18500+8500+40900+3200)-(278600+8600+14300+1750+26200) = $       80,500
Q-2) Cash surplus/deficit = 72400-7550-19500-4200-9100+400-8450-6600-5000-5500 = $          6,900
Q-3) Solvency ratio for the current year = Net worth/Total assets = 80500/(1500+2350+335000+18500+8500+40900+3200)= 19.64%
As the solvency ratio was 42% the previous year,
*there is a drastic decline from the previous year.